Devyn and Sandra's first two children are boys with their mother's big brown eyes. Their third...

  1. Devyn and Sandra's first two children are boys with their mother's big brown eyes. Their third child is a daughter who has her father Devyn's blue eyes. Blue eyes are recessive to brown. What is the probability that Devyn and Sandra's 4th child will also have blue eyes? Please provide your answer as a decimal to two decimal places (e.g. 0.75). Do not forget to round up or down, as appropriate.

    What is the probability that Devyn and Sandra's 4th child will be another blue-eyed girl? Please provide your answer as a decimal to two decimal places (e.g. 0.75). Do not forget to round up or down, as appropriate.
Answer #1

Give blue (b) are recessive to brown (B).
Sandra= Brown eyes (Bb) (Since one of their child exhibit recessive phenotype, the dominant parent should be heterozygous)
Devan= blue eyes (bb)
Probability of their 4th child to have blue eyes= Bb x bb = 1/2 bb = 0.50

b) Probability of their 4th child to be blue eyes girl :
Probability of having blue eyed child = 1/2
probability of having a girl child= 1/2
So, total probability of having a blue eyed girl child = 1/2 x 1/2 = 1/4 = 0.25.